OpenStudy (nuccioreggie):

A proportional relationship between the number of pounds of potatoes (x) and the price in dollars (y) is graphed, and the ordered pair (4, 3) is on the graphed line. Part A: What is the price of 1 pound of potatoes? Show your work, including the proportion you used to determine the price. (8 points) Part B: What does the ordered pair (8, 6) on the graph represent? Explain in words. (2 points)

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (michele_laino):

the unit price is given by the subsequent ratio: 3/4 =...?

OpenStudy (michele_laino):

part B: 6 is the price for 8 pounds of potatoes

OpenStudy (nuccioreggie):

Ok but @Michele_Laino. It say i have to show my work and the proportion i used to determine the price

OpenStudy (michele_laino):



OpenStudy (michele_laino):

part A. I call with p the price of 1 pound of potatoes, so I can write this proportion: \[\Large 3:4 = p:1\]

OpenStudy (nuccioreggie):

Im talking about in words @Michele_Laino

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (michele_laino):

here is my answer: "three over four is equals to p over one"

OpenStudy (michele_laino):

@nuccioreggie

OpenStudy (nuccioreggie):

O ok and what is B @Michele_Laino

OpenStudy (nuccioreggie):

???

OpenStudy (michele_laino):

Part B: "6 is the price of 8 pounds of potatoes"

OpenStudy (nuccioreggie):

How did we get that / know that

OpenStudy (nuccioreggie):

@Michele_Laino. In part A how did you come up with that answer

OpenStudy (michele_laino):

since if 4 pounds cost $3 than 1 pound costs p, namely those ratios 4/3 and 1/p have to be equal each other: \[\frac{4}{3} = \frac{1}{p}\]

OpenStudy (michele_laino):

or, taking the inverse of those ratios, we can write: \[\frac{3}{4} = \frac{p}{1}\]

OpenStudy (nuccioreggie):

Ok thank you i have 3 more ok